The temperature in a refrigerator coil is 267K and that in the condenser coil is 295K. Assuming that the machine operates on the reversed Carnot cycle, calculate :

(i) C.O.P.(ref.)

(ii) The refrigerating effect per kW of input work.

(iii) The heat rejected to the condenser.




1
Expert's answer
2021-07-20T14:59:37-0400
"Cop=\frac{T_2}{T_1-T_2}=\frac{295}{295-267}=\frac{295}{28}=10.53"

(2) COP ="\frac{refrigerating effect}{work input}"

Work input ="\\\frac{3.6\times10^6}{10.53}=341.88KJ" (3)

Heat rejected to the condensor =COP"\times" work done

"10.28\times341.88=3514.52" kJ